# MagicDistillation: Weak-to-Strong Video Distillation for Large-Scale Portrait Few-Step Synthesis

### v1

@@ 4.2 Ablation Study
In this subsection, we perform ablation studies on the hyperparameter $\alpha_{\textrm{weak}}$ and the loss $\mathcal{L}_{\textrm{reg}}$. We present the results of the ablation experiments on $\alpha_{\textrm{weak}}$ and $\mathcal{L}_{\textrm{reg}}$ in Fig. 5. When $\alpha_{\textrm{weak}}$ is set to 0.25, it achieves the best performance in the i2v subject and dynamic degree metrics. Furthermore, as $\alpha_{\textrm{weak}}$ gradually increases, both aesthetic quality and imaging quality improve monotonically, indicating that increasing $\alpha_{\textrm{weak}}$ enhances the visual quality of the synthesized videos. Conversely, when $\alpha_{\textrm{weak}}$ increases from 0.5 to 0.75, the dynamic degree drops sharply, suggesting that an excessively high $\alpha_{\textrm{weak}}$ negatively impacts the motion dynamics of the synthesized videos. Overall, the average score reaches its optimum when $\alpha_{\textrm{weak}}=0.25$. Thus, we adopt a default configuration of $\alpha_{\textrm{weak}}=0.25$ and $\alpha_{\textrm{strong}}=1$. Furthermore, the observation that ground truth supervision substantially enhances both aesthetic quality and imaging quality while adversely affecting the dynamic degree aligns closely with the findings reported in the above main results.
@@ 4.3 Visualization
We provide a visual comparison of Euler, LCM, DMD2, and W2SVD in Fig. 6. More comparison videos can be found in project page. As illustrated in Fig. 6, both LCM and Euler sampling produce blurry results in the 1/4-step scenarios, making it difficult to synthesize high-quality talking videos. In contrast, DMD2 synthesizes relatively clearer videos in the 1/4-step scenarios but still suffers from uncontrollable synthesis issues. For example, in the 4-step scenario, the animated portrait videos synthesized by DMD2 exhibit significant changes in the subject’s facial features. In comparison, W2SVD successfully synthesizes high-fidelity videos in both the 1-step and 4-step scenarios, achieving superior visual quality, subject consistency, and motion dynamics.
@@ 5 Conclusion
In this paper, we introduce a more stable, efficient, and effective step distillation algorithm, termed W2SVD, designed for accelerated sampling of large-scale VDM. Our work centers around the portrait video synthesis task, leveraging the 13b+ HunyuanVideo model alongside LoRA to facilitate weak-to-strong distribution matching. This approach allows for a more reliable estimation of $\mathcal{D}_{\textrm{KL}}(p_{\textrm{fake}}\|p_{\textrm{real}})$. Furthermore, we propose incorporating ground truth supervision to achieve a useful trade-off between visual quality and motion dynamics. Extensive experimental results demonstrate that W2SVD surpasses DMD2 and even the 28-step teacher model in nearly all metrics.
@@
We present W2SVD, a method designed for realizing more efficient and better step distillation on large scale VDMs. The dataset utilized by W2SVD is sourced from animated video clips and licensed public portrait videos available online. This dataset undergoes meticulous filtering through both manual review and large language model (LLM)-based processes to ensure that all reference images adhere to ethical standards. Furthermore, we place a strong emphasis on the responsible and ethical deployment of this technology, striving to maximize its societal benefits while diligently addressing and mitigating any potential risks.
@@ Appendix A Benchmarks
In portrait video synthesis, the primary metrics for evaluating video quality are FID (Heusel et al., 2017) and FVD (Unterthiner et al., 2019). Following EMO (Sun et al., 2023) and Hallo (Cui et al., 2024a), we randomly select 100 clips from each of the VFHQ (Xie et al., 2022), HDTF (Zhang et al., 2021), and Celeb-V (Zhu et al., 2022) datasets. Each clip consists of 129 frames, with the resolution and frame rate standardized to 540$\times$960 and 24 FPS, respectively. We then compute the FID and FVD scores by comparing the synthesized videos with the original videos.
@@
Evaluating videos solely based on FVD and FID has clear limitations in terms of metric coverage. However, most of the existing video benchmarks, such as Chronomagic-Bench (Yuan et al., 2024) and T2V-CompBench (Sun et al., 2024), are primarily designed for T2V generation, with only a limited number tailored for I2V synthesis. To address this gap, we adopt the I2V-VBench construction framework from VBench++ (Huang et al., 2024b). Specifically, we begin by collecting 58 high-quality widescreen images, and use InternVL-26B to generate prompts for these images, and then synthesize five videos per sample (resulting in a total of 290 videos) to minimize testing errors. The synthesized videos are subsequently evaluated using VBench. Note that the key evaluation metrics include: i2v subject, subject consistency, motion smoothness, dynamic degree, aesthetic quality, imaging quality, and temporal flickering.
@@ Appendix B Hyperparameter Settings
The LCM algorithm utilized in our training process largely follows the methodology outlined in FastVideo (Hao-AI-Lab, 2025). Specifically, we set the number of timesteps for training to 28, with a batch size of 16 (equivalent to performing gradient accumulation four times). The learning rate is fixed at 1e-6 throughout the training process, and loss matching is conducted at $t=0$ using the Huber loss function. It is important to note that, due to GPU memory constraints, we adopt the same approach as FastVideo by not employing EMA. Instead, we treat the target model and the teacher model as identical during LCM training. For evaluation, we use the checkpoint obtained after completing 20k iterations.
@@
W2SVD.
The majority of the experimental setup for W2SVD aligns with that of vanilla DMD/DMD2. Unless specified otherwise in the ablation experiments, we set $\alpha_{\textrm{strong}}$ and $\alpha_{\textrm{weak}}$ to 1 and 0.25, respectively, as the default configuration. For the 1-step generator, the pre-trained model is trained with $\alpha_{\textrm{weak}}=0.25$ and optimized under the regularization loss $\mathcal{L}_{\textrm{reg}}$ scenario using W2SVD. As W2SVD significantly mitigates the problem of training collapse, we select checkpoints obtained after the 600th iteration for the 4-step scenario and after the 1500th iteration for the 1-step scenario for evaluation.
@@
$\begin{split}&\mathop{\arg\min}_{\phi}\ \mathbb{E}_{t,(\epsilon,\epsilon^{\prime})\sim\mathcal{N}(0,\mathbf{I})}\left[\|\epsilon^{\prime}-G_{\phi}(\epsilon)-\mathbf{v}_{\Theta}^{\textrm{real}}(\mathbf{x}_{t},t)\|_{2}^{2}\right]\\ &\implies\mathop{\arg\min}_{\phi}\ \mathbb{E}_{t,(\epsilon,\epsilon^{\prime})\sim\mathcal{N}(0,\mathbf{I})}\left[\|\epsilon^{\prime}-G_{\phi}(\epsilon)-\mathbf{v}_{\Theta}^{\textrm{pre-train}}(\mathbf{x}_{t},t)-\alpha_{\textrm{weak}}\zeta(\mathbf{x}_{t},t)\|_{2}^{2}\right]\\ &\implies\mathop{\arg\min}_{\phi}\ \mathbb{E}_{t,(\epsilon,\epsilon^{\prime})\sim\mathcal{N}(0,\mathbf{I})}\left[\|G_{\phi}^{*}(\epsilon)-G_{\phi}(\epsilon)-\alpha_{\textrm{weak}}\zeta(\mathbf{x}_{t},t)\|_{2}^{2}\right]\\ &\implies\mathop{\arg\min}_{\phi}\ \mathbb{E}_{t,(\epsilon,\epsilon^{\prime})\sim\mathcal{N}(0,\mathbf{I})}\left[\left\|\left(\frac{\alpha_{\textrm{strong}}-\alpha_{\textrm{weak}}}{\alpha_{\textrm{strong}}}\right)[G_{\phi}^{*}(\epsilon)-G_{\phi}(\epsilon)]\right\|\right]\\ &\implies\mathop{\arg\min}_{\phi}\ \mathbb{E}_{t,(\epsilon,\epsilon^{\prime})\sim\mathcal{N}(0,\mathbf{I})}\left(\frac{\alpha_{\textrm{strong}}-\alpha_{\textrm{weak}}}{\alpha_{\textrm{strong}}}\right)^{2}\left[\left\|\epsilon^{\prime}-\mathbf{v}^{\textrm{pre-train}}_{\Theta}(\sigma_{t}\epsilon^{\prime}+(1-\sigma_{t})G_{\phi}(\epsilon),t)-G_{\phi}(\epsilon)\right\|_{2}^{2}\right]\\ &\implies\mathop{\arg\min}_{\phi}\ \mathbb{E}_{t,(\epsilon,\epsilon^{\prime})\sim\mathcal{N}(0,\mathbf{I})}\left(\frac{\alpha_{\textrm{strong}}-\alpha_{\textrm{weak}}}{\alpha_{\textrm{strong}}}\right)^{2}\left[\left\|\mathbf{v}_{\theta}^{\textrm{fake}}(\sigma_{t}\epsilon^{\prime}+(1-\sigma_{t})G_{\phi}(\epsilon),t)-\mathbf{v}^{\textrm{pre-train}}_{\Theta}(\sigma_{t}\epsilon^{\prime}+(1-\sigma_{t})G_{\phi}(\epsilon),t)\right\|_{2}^{2}\right].\\ \end{split}$ (10)
@@
Appendix D Experimental Result of TI2V-HunyuanVideo
Herein, we present the performance of W2SVD on TI2V-HunyuanVideo. Following the identical experimental setup employed for W2SVD in I2V-HunyuanVideo, we enhanced computational resources to 8 nodes with 40 NVIDIA H100 GPUs, aiming to achieve superior performance. As demonstrated in Fig. 7 and Table 5 containing visualizations and quantitative evaluations, the distilled 4-step generator through W2SVD exhibits enhanced synthesis fidelity and improved motion consistency in TI2V portrait video generation tasks. These empirical results comprehensively validate the technical advantages of W2SVD, particularly highlighting its capability to preserve temporal coherence while maintaining high visual quality in compressed-step video synthesis scenarios.
Table 5: Quantitative results comparison using TI2V-HunyuanVideo (finetuned on portrait video synthesis task) on our customized VBench. All videos are 540$\times$960$\times$129 and are saved at 24fps. Each sample follow VBench (Huang et al., 2024a) to synthesize 5 videos to avoid errors.
@@ Appendix D Experimental Result of TI2V-HunyuanVideo
W2SVD w/ reg. loss (ours) 0.9807 0.9777 0.9951 0.0620 0.5979 0.7110 0.9974 0.7602
Figure 7: Visualization of W2SVD and the standard Euler sampling on the TI2V-HunyuanVideo model. In the 4-step scenario, W2SVD visibly outperforms Euler in terms of visual quality and facial motion dynamics.
Appendix E Pseudo Code of W2SVD
In order to give the researchers a better understanding of W2SVD, we have placed the pseudo-code of this method in Algorithm 1.
Algorithm 1 Weak-to-Strong Video Distillation (W2SVD) Training Procedure
@@ Appendix E Pseudo Code of W2SVD
// Initialize the few-step generator and the real data distribution estimator $\mathbf{v}_{\Theta}^{\textrm{real}}$, the fake data distribution estimator $\mathbf{v}_{\theta}^{\textrm{fake}}$
$G_{\phi}\leftarrow\text{copyWeights}(\mathbf{v}_{\Theta}^{\textrm{pre-train}}),$ $\mathbf{v}^{\textrm{real}}_{\Theta}\leftarrow\alpha_{\textrm{weak}}\zeta+\mathbf{v}_{\Theta}^{\textrm{pre-train}},$ $\mathbf{v}^{\textrm{fake}}_{\theta}\leftarrow\alpha_{\textrm{strong}}\zeta+\mathbf{v}_{\Theta}^{\textrm{pre-train}}$
while train do
2       // Generate videos Sample batch $\epsilon\sim\mathcal{N}(0,\mathbf{I})^{B}$, the ground truth video $\mathbf{x}^{\textrm{gt}}\sim\mathcal{D}$ and $t\sim[t_{1},\cdots,t_{N}]$   // Prepare for W2SVD Get the noisy video $\mathbf{x}_{t}=(t/1000)\epsilon+((1-t)/1000)\mathbf{x}^{\textrm{gt}}$   // Prepare for W2SVD Get the synthesized noisy video $\mathbf{\tilde{x}}_{t}=(t/1000)\epsilon+((1-t)/1000)[\mathbf{x}_{t}-tG_{\phi}(\mathbf{x}_{t},t)]$   // Prepare for W2SVD $\mathcal{L}_{\text{DMD}}\leftarrow\text{distributionMatchingLoss}(\mathbf{v}^{\textrm{real}}_{\Theta},\mathbf{v}^{\textrm{fake}}_{\theta},\mathbf{\tilde{x}}_{t})$ // Eq 2 $\mathcal{L}_{\text{reg}}\leftarrow\text{distributionMatchingLoss}(\mathbf{x}^{\textrm{gt}},\mathbf{v}^{\textrm{fake}}_{\theta},\mathbf{\tilde{x}}_{t})$ // Eq 7 $\mathcal{L}_{\text{gen}}\leftarrow\text{generatorLoss}(D_{\xi},G_{\phi},\epsilon,\mathbf{x}^{\textrm{gt}})$ // Eq 4 $\mathcal{L}_{G_{\phi}}\leftarrow\mathcal{L}_{\text{DMD}}+\mathcal{L}_{\text{reg}}+\mathcal{L}_{\text{dis}}$ // Get the overall loss $G_{\phi}\leftarrow\text{update}(G_{\phi},\mathcal{L}_{G_{\phi}})$   // Update the few-step generator Sample time step $t\sim\mathcal{U}[0,1]$ and the Gaussian noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})^{B}$ $\mathbf{x}^{\textrm{fake}}_{0}\leftarrow\mathbf{x}_{t}-tG_{\phi}(\mathbf{x}_{t},t)$ $\mathbf{x}^{\textrm{fake}}_{t}\leftarrow(t/1000)\epsilon+((1-t)/1000)\mathbf{x}^{\textrm{fake}}_{0}$ $\mathcal{L}_{\text{diffusion}}\leftarrow\text{diffusionLoss}(\mathbf{v}^{\text{fake}}_{\theta}(\mathbf{x}^{\textrm{fake}}_{t},t),\text{stopgrad}(\mathbf{x}^{\textrm{fake}}_{0}))$ // Eq 3 $\mathcal{L}_{\text{dis}}\leftarrow\text{discriminatorLoss}(D_{\xi},G_{\phi},\epsilon)$ // Eq 4 $\zeta\leftarrow\text{update}(\zeta,\mathcal{L}_{\text{diffusion}}+\mathcal{L}_{\text{dis}})$  // Update the low-rank branch

### v2

@@ 4.2 Ablation Study
In this subsection, we perform ablation studies on the hyperparameter $\alpha_{\textrm{weak}}$ and the loss $\mathcal{L}_{\textrm{reg}}$. We present the results of the ablation experiments on $\alpha_{\textrm{weak}}$ and $\mathcal{L}_{\textrm{reg}}$ in Fig. 4. When $\alpha_{\textrm{weak}}$ is set to 0.25, it achieves the best performance in the i2v subject and dynamic degree metrics. Furthermore, as $\alpha_{\textrm{weak}}$ gradually increases, both aesthetic quality and imaging quality improve monotonically, indicating that increasing $\alpha_{\textrm{weak}}$ enhances the visual quality of the synthesized videos. Conversely, when $\alpha_{\textrm{weak}}$ increases from 0.5 to 0.75, the dynamic degree drops sharply, suggesting that an excessively high $\alpha_{\textrm{weak}}$ negatively impacts the motion dynamics of the synthesized videos. Overall, the average score reaches its optimum when $\alpha_{\textrm{weak}}=0.25$. Thus, we adopt a default configuration of $\alpha_{\textrm{weak}}=0.25$ and $\alpha_{\textrm{strong}}=1$. Furthermore, the observation that ground truth supervision substantially enhances both aesthetic quality and imaging quality while adversely affecting the dynamic degree. Based on MagicDistillation’s performance in Tables 1 and 2, we believe that the detrimental effects of $\mathcal{L}_{\textrm{reg}}$ can be avoided as long as the training data is mixed with general data exhibiting high motion dynamics.
@@ 4.3 Visualization
We provide a visual comparison of Euler, LCM, DMD2, and MagicDistillation in Figs. 6 and 11. More results can be found in project page. As illustrated in Figs. 6 and 11, both LCM and Euler sampling produce blurry results in the 1/4-step scenarios, making it difficult to synthesize high-quality talking videos. In contrast, DMD2 synthesizes relatively clearer videos in the 1/4-step scenarios but still suffers from uncontrollable synthesis issues. For example, in the 4-step scenario, the animated portrait videos synthesized by DMD2 exhibit significant changes in the subject’s facial features. In comparison, MagicDistillation successfully synthesizes high-fidelity videos in both the 1-step and 4-step scenarios, achieving superior visual quality, subject consistency, and motion dynamics.
@@ 5 Conclusion
In this paper, we introduce a more stable, efficient, and effective step distillation algorithm, termed MagicDistillation, designed for accelerated sampling of large-scale VDM. Our research centers on addressing the substantial inference overhead posed by large-scale VDMs while simultaneously enhancing the performance of few-step generators in the realm of portrait video synthesis. We leveraging the 13B Magic141 model alongside LoRA to facilitate weak-to-strong distribution matching. This approach allows for a more reliable estimation of $\mathcal{D}_{\textrm{KL}}(p_{\textrm{fake}}\|p_{\textrm{real}})$. Furthermore, we propose incorporating ground truth supervision to achieve superior visual quality. Extensive experimental results demonstrate that MagicDistillation surpasses DMD, DMD2, LCM, the 28/50-step Magic141, the 50-step WanX-I2V and the 50-step HunyuanVideo-I2V in nearly all metrics.
@@
We present MagicDistillation, a method designed for realizing more efficient and better step distillation on large scale VDMs. The dataset utilized by MagicDistillation is sourced from animated video clips and licensed public portrait videos available online. This dataset undergoes meticulous filtering through both manual review and large language model (LLM)-based processes to ensure that all reference images adhere to ethical standards. Furthermore, we place a strong emphasis on the responsible and ethical deployment of this technology, striving to maximize its societal benefits while diligently addressing and mitigating any potential risks.
@@ Appendix A Benchmarks
In portrait video synthesis, the primary metrics for evaluating video quality are FID [8] and FVD [33]. Following EMO [30] and Hallo [4], we randomly select 100 clips from each of the VFHQ [37], HDTF [49], and Celeb-V [51] datasets. Each clip consists of 129 frames, with the resolution and frame rate standardized to 540$\times$960 and 24 FPS, respectively. We then compute the FID and FVD scores by comparing the synthesized videos with the original videos.
@@
Evaluating videos solely based on FVD and FID has clear limitations in terms of metric coverage. However, most of the existing video benchmarks, such as Chronomagic-Bench [44] and T2V-CompBench [29], are primarily designed for T2V generation, with only a limited number tailored for I2V synthesis. To address this gap, we adopt the I2V-VBench construction framework from VBench++ [11]. Specifically, we begin by collecting 58 high-quality widescreen images, and use InternVL-26B to generate prompts for these images, and then synthesize five videos per sample (resulting in a total of 290 videos) to minimize testing errors. The synthesized videos are subsequently evaluated using VBench. Note that the key evaluation metrics include: i2v subject, subject consistency, motion smoothness, dynamic degree, aesthetic quality, imaging quality, and temporal flickering.
General I2V-VBench.
The original VBench [10] was designed solely to evaluate the generation performance of T2V-generators and lacked the capability to assess the performance of I2V-generators. Thus, I2V-VBench [11] is an extension built upon the vanilla VBench framework. This enhancement incorporates over 1118 additional text-image pairs, enabling a comprehensive evaluation of I2V-generators across multiple dimensions, including visual quality, motion dynamics, and subject consistency in the synthesized videos. For the evaluation, MagicDistillation is assessed using a 16:9 resolution ratio, and the performance metrics of the comparison methods are sourced directly from the official leaderboard available at https://huggingface.co/spaces/Vchitect/VBench_Leaderboard.
@@ Appendix B Hyperparameter Settings
The LCM algorithm utilized in our training process largely follows the methodology outlined in FastVideo [7]. Specifically, we set the number of timesteps for training to 28, with a batch size of 16 (equivalent to performing gradient accumulation four times). The learning rate is fixed at 1e-6 throughout the training process, and loss matching is conducted at $t=0$ using the Huber loss function. It is important to note that, due to GPU memory constraints, we adopt the same approach as FastVideo by not employing EMA. Instead, we treat the target model and the teacher model as identical during LCM training. For evaluation, we use the checkpoint obtained after completing 20k iterations.
@@
MagicDistillation.
The majority of the experimental setup for MagicDistillation aligns with that of vanilla DMD/DMD2. Unless specified otherwise in the ablation experiments, we set $\alpha_{\textrm{strong}}$ and $\alpha_{\textrm{weak}}$ to 1 and 0.25, respectively, as the default configuration. For the 1-step generator, the pre-trained model is trained with $\alpha_{\textrm{weak}}=0.25$ and optimized under the regularization loss $\mathcal{L}_{\textrm{reg}}$ scenario using MagicDistillation. As MagicDistillation significantly mitigates the problem of training collapse, we select checkpoints obtained after the 600th iteration for the 4-step scenario and after the 1500th iteration for the 1-step scenario for evaluation.
@@
$\begin{split}&\mathop{\arg\min}_{\phi}\ \mathbb{E}_{t,(\epsilon,\epsilon^{\prime})\sim\mathcal{N}(0,\mathbf{I})}\left[\|\epsilon^{\prime}-G_{\phi}(\epsilon)-\mathbf{v}_{\Theta}^{\textrm{real}}(\mathbf{x}_{t},t)\|_{2}^{2}\right]\\ &\implies\mathop{\arg\min}_{\phi}\ \mathbb{E}_{t,(\epsilon,\epsilon^{\prime})\sim\mathcal{N}(0,\mathbf{I})}\left[\|\epsilon^{\prime}-G_{\phi}(\epsilon)-\mathbf{v}_{\Theta}^{\textrm{pre-train}}(\mathbf{x}_{t},t)-\alpha_{\textrm{weak}}\zeta(\mathbf{x}_{t},t)\|_{2}^{2}\right]\\ &\implies\mathop{\arg\min}_{\phi}\ \mathbb{E}_{t,(\epsilon,\epsilon^{\prime})\sim\mathcal{N}(0,\mathbf{I})}\left[\|G_{\phi}^{*}(\epsilon)-G_{\phi}(\epsilon)-\alpha_{\textrm{weak}}\zeta(\mathbf{x}_{t},t)\|_{2}^{2}\right]\\ &\implies\mathop{\arg\min}_{\phi}\ \mathbb{E}_{t,(\epsilon,\epsilon^{\prime})\sim\mathcal{N}(0,\mathbf{I})}\left[\left\|\left(\frac{\alpha_{\textrm{strong}}-\alpha_{\textrm{weak}}}{\alpha_{\textrm{strong}}}\right)[G_{\phi}^{*}(\epsilon)-G_{\phi}(\epsilon)]\right\|\right]\\ &\implies\mathop{\arg\min}_{\phi}\ \mathbb{E}_{t,(\epsilon,\epsilon^{\prime})\sim\mathcal{N}(0,\mathbf{I})}\left(\frac{\alpha_{\textrm{strong}}-\alpha_{\textrm{weak}}}{\alpha_{\textrm{strong}}}\right)^{2}\left[\left\|\epsilon^{\prime}-\mathbf{v}^{\textrm{pre-train}}_{\Theta}(\sigma_{t}\epsilon^{\prime}+(1-\sigma_{t})G_{\phi}(\epsilon),t)-G_{\phi}(\epsilon)\right\|_{2}^{2}\right]\\ &\implies\mathop{\arg\min}_{\phi}\ \mathbb{E}_{t,(\epsilon,\epsilon^{\prime})\sim\mathcal{N}(0,\mathbf{I})}\left(\frac{\alpha_{\textrm{strong}}-\alpha_{\textrm{weak}}}{\alpha_{\textrm{strong}}}\right)^{2}\\ &\left[\left\|\mathbf{v}_{\theta}^{\textrm{fake}}(\sigma_{t}\epsilon^{\prime}+(1-\sigma_{t})G_{\phi}(\epsilon),t)-\mathbf{v}^{\textrm{pre-train}}_{\Theta}(\sigma_{t}\epsilon^{\prime}+(1-\sigma_{t})G_{\phi}(\epsilon),t)\right\|_{2}^{2}\right].\\ \end{split}$ (10)
@@
Appendix D Why We Choice DMD?
Figure 7: Comparison of LCM, DMD and SiD training paradigms.
Most existing methods for step distillation in VDMs are built upon the latent consistency model (LCM) paradigm [26]. However, as illustrated in Fig. 7, the LCM framework necessitates the incorporation of the exponential moving average (EMA) mechanism, resulting in the requirement to load three models’ parameters during training. Unfortunately, through our exploratory studies, we observed that even with the use of 16 GPUs equipped with 80GB and the zero redundancy optimizer (ZeRO) stage 3 CPU offload technique [22], the out-of-memory (OOM) issue persists and cannot be resolved.
Consequently, we turn our attention to an EMA-independent step distillation method known as score identity distillation (SiD) [50] and distribution matching distillation (DMD) [42, 41]. Compared to DMD, SiD necessitates calculating the gradients of both the real model and the fake model during the update of the few-step generator, resulting in significantly lower computational efficiency. Consequently, we ultimately chose DMD, which is expected to enable the development of the few-step generator, which theoretically achieves both efficiency and effectiveness. In detail, DMD aims to estimate both the real data distribution and the fake data distribution using two distinct models with non-overlapping parameters. The generator’s parameters are then updated using the score distillation sampling (SDS) algorithm [21], leveraging information from both distributions. It is clear that the training process of vanilla DMD similarly requires three models, which can result in the same OOM issue as encountered with LCM. To address this issue without compromising the generator’s final performance, we propose adding the low-rank adaptation (LoRA) on the real model (i.e., estimate real data distribution) to implement the fake model (i.e., estimate fake data distribution), ensuring efficient training and avoiding the OOM issue.
Appendix E Experimental Result of Magic141-TI2V
Herein, we present the performance of MagicDistillation on TI2V-HunyuanVideo. Following the identical experimental setup employed for MagicDistillation in I2V-HunyuanVideo, we enhanced computational resources to 8 nodes with 40 NVIDIA H100 GPUs, aiming to achieve superior performance. As demonstrated in Fig. 8 and Table 7 containing visualizations and quantitative evaluations, the distilled 4-step generator through MagicDistillation exhibits enhanced synthesis fidelity and improved motion consistency in TI2V portrait video generation tasks. These empirical results comprehensively validate the technical advantages of MagicDistillation, particularly highlighting its capability to preserve temporal coherence while maintaining high visual quality in compressed-step video synthesis scenarios.
Table 7: TI2V quantitative results comparison using Magic141 (trained on both portrait and general video datasets) on our customized VBench. All videos are 540$\times$960$\times$129 and are saved at 24fps. Each sample follow VBench [10] to synthesize 5 videos to avoid errors.
@@ Appendix E Experimental Result of Magic141-TI2V
MagicDistillation w/ reg. loss (ours) 0.9807 0.9777 0.9951 0.0620 0.5979 0.7110 0.9974 0.7602
Figure 8: Visualization of MagicDistillation and the standard Euler sampling on the TI2V-HunyuanVideo model. In the 4-step scenario, MagicDistillation visibly outperforms Euler in terms of visual quality and facial motion dynamics.
Appendix F Pseudo Code of MagicDistillation
In order to give the researchers a better understanding of MagicDistillation, we have placed the pseudo-code of this method in Algorithm 1.
@@ Appendix F Pseudo Code of MagicDistillation
1 // Initialize the few-step generator and the real data distribution estimator $\mathbf{v}_{\Theta}^{\textrm{real}}$, the fake data distribution estimator $\mathbf{v}_{\theta}^{\textrm{fake}}$
2 $G_{\phi}\leftarrow\text{copyWeights}(\mathbf{v}_{\Theta}^{\textrm{pre-train}}),$ $\mathbf{v}^{\textrm{real}}_{\Theta}\leftarrow\alpha_{\textrm{weak}}\zeta+\mathbf{v}_{\Theta}^{\textrm{pre-train}},$ $\mathbf{v}^{\textrm{fake}}_{\theta}\leftarrow\alpha_{\textrm{strong}}\zeta+\mathbf{v}_{\Theta}^{\textrm{pre-train}}$
3 while train do
4       // Generate videos
5       Sample batch $\epsilon\sim\mathcal{N}(0,\mathbf{I})^{B}$, the ground truth video $\mathbf{x}^{\textrm{gt}}\sim\mathcal{D}$ and $t\sim[t_{1},\cdots,t_{N}]$   // Prepare for MagicDistillation
6       Get the noisy video $\mathbf{x}_{t}=(t/1000)\epsilon+((1-t)/1000)\mathbf{x}^{\textrm{gt}}$   // Prepare for MagicDistillation
7       Get the synthesized noisy video $\mathbf{\tilde{x}}_{t}=(t/1000)\epsilon+((1-t)/1000)[\mathbf{x}_{t}-tG_{\phi}(\mathbf{x}_{t},t)]$   // Prepare for MagicDistillation
8       $\mathcal{L}_{\text{DMD}}\leftarrow\text{distributionMatchingLoss}(\mathbf{v}^{\textrm{real}}_{\Theta},\mathbf{v}^{\textrm{fake}}_{\theta},\mathbf{\tilde{x}}_{t})$ // Eq 2
9       $\mathcal{L}_{\text{reg}}\leftarrow\text{distributionMatchingLoss}(\mathbf{x}^{\textrm{gt}},\mathbf{v}^{\textrm{fake}}_{\theta},\mathbf{\tilde{x}}_{t})$ // Eq 7
10       $\mathcal{L}_{\text{gen}}\leftarrow\text{generatorLoss}(D_{\xi},G_{\phi},\epsilon,\mathbf{x}^{\textrm{gt}})$ // Eq 4
11       $\mathcal{L}_{G_{\phi}}\leftarrow\mathcal{L}_{\text{DMD}}+\mathcal{L}_{\text{reg}}+\mathcal{L}_{\text{dis}}$ // Get the overall loss
12       $G_{\phi}\leftarrow\text{update}(G_{\phi},\mathcal{L}_{G_{\phi}})$   // Update the few-step generator
13       Sample time step $t\sim\mathcal{U}[0,1]$ and the Gaussian noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})^{B}$
14      $\mathbf{x}^{\textrm{fake}}_{0}\leftarrow\mathbf{x}_{t}-tG_{\phi}(\mathbf{x}_{t},t)$
15      $\mathbf{x}^{\textrm{fake}}_{t}\leftarrow(t/1000)\epsilon+((1-t)/1000)\mathbf{x}^{\textrm{fake}}_{0}$
16      $\mathcal{L}_{\text{diffusion}}\leftarrow\text{diffusionLoss}(\mathbf{v}^{\text{fake}}_{\theta}(\mathbf{x}^{\textrm{fake}}_{t},t),\text{stopgrad}(\mathbf{x}^{\textrm{fake}}_{0}))$ // Eq 3
17       $\mathcal{L}_{\text{dis}}\leftarrow\text{discriminatorLoss}(D_{\xi},G_{\phi},\epsilon)$ // Eq 4
18
19      $\zeta\leftarrow\text{update}(\zeta,\mathcal{L}_{\text{diffusion}}+\mathcal{L}_{\text{dis}})$  // Update the low-rank branch
20